Volcano Hot Wings has budgeted the following costs for a month in which 12,000 wings will be cooked and sold.

Wings, breading, and sauce $ 4,900
Direct labor (Variable) 3,500
Rent 1,100
Depreciation 900
Other fixed costs 400


Each wing sells for $1.60 each. How much would Volcano Hot Wings’ profit increase if 100 more wings were sold?
Answer #1
Amount $
Incremental revenue     160.00 =100*1.6
Less: Increase Variable costs
Direct labor        29.17 =3500/12000*100
Wings, breading, and sauce        40.83 =4900/1200*100
Increase in profit        90.00